OpenStudy (anonymous):

solving equations by elimination!! -9x + 3y = 27 10x - 3y = -30

OpenStudy (goformit100):

solve it you can do it. I Trust you can do it.. It's Easy

OpenStudy (goformit100):

@pinklove31419 you ca do it

OpenStudy (anonymous):

Add them toghether together. So -9x +10x = x 3y+(-3y)=0 so x =27-30

OpenStudy (anonymous):

KNOW you now 'x' you can work out 'y'

OpenStudy (anonymous):

-9x + 3y = 27 10x - 3y = -30 --------------add x = - 3 (as you can see, the y's cancel each other out) now substitute -3 in for x in either of the original equations.... -9(-3) + 3y = 27 or 10(-3) - 3y = -30 both of these equations, when worked out, will give you the same answer for y.